OpenStudy (anonymous):

Suppose oil spills from a ruptured tanker and spreads in a circular pattern. If the radius of the oil spill increases at a constant rate of 3m/s, how fast is the area of the spill increasing when the radius is 40m? Answer in m^2/s

OpenStudy (anonymous):

@thomaster @robtobey @phi

OpenStudy (phi):

 \[ \frac{d}{dt}\left(A= \pi r^2 \right)\\ \frac{d}{dt}A= \pi \frac{d}{dt}r^2\\ \frac{dA}{dt}= 2 \pi r \frac{dr}{dt} \] radius of the oil spill increases at a constant rate of 3m/s means dr/dt = 3 m/s when the radius is 40m? i.e. r= 40 m find dA/dt

OpenStudy (anonymous):

i dont get how to find it.

OpenStudy (phi):

dA/dt is the change in area with respect to time. dA/dt is how fast is the area of the spill increasing on the right side of the equation, you know all of the values.

OpenStudy (anonymous):

the right side would be \[6\pi r\]?

OpenStudy (phi):

how fast is the area of the spill increasing when the radius is 40m?

OpenStudy (anonymous):

i dont know. I dont get this

OpenStudy (phi):

the idea is not too tricky, but sometimes things don't "click" but in this case, r is the radius. they want dA/dt when r= 40 m, and dr/dt= 3 m/s

OpenStudy (anonymous):

so when i plug in dA/dt and dr/dt into the equation am i supposed to put it as 40 and 3 or as something else

OpenStudy (phi):

You have an equation \[ \frac{dA}{dt}= 2 \pi r \frac{dr}{dt} \] you know r=40 and dr/dt =3 replace r with 40 and dr/dt with 3 simplify to find dA/dt

OpenStudy (anonymous):

so its 240pi

OpenStudy (phi):

or about 754 m^2/sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how would you do this for any other situation

OpenStudy (phi):

the idea is if you know (in a very simple example) y = x and x changes, then y changes. we can take the derivative with respect to time: dy/dt = dx/dt that says that if the change in x is dx/dt (think of this as a rate) then the change in y will be the same rate. This should make sense, because y=x and if x is changing at 1 step per second, y will be changing at 1 step per second.

OpenStudy (phi):

if y= 2x if x changes at 1 step per second, y will go up by 2 steps per second take the derivative with respect to time dy/dt = 2 dx/dt that says the change in y is twice as big as the change in x. in other words, if x goes up by 1 step every second, y will be going up at 2 steps every second.

OpenStudy (anonymous):

So with a problem like this: The length of a rectangle is increasing at a rate of 8cm/s and its width is increasing at a rate of 5cm/s. When the length is 30cm and the width is 25cm, how fast is the area of the rectangle increasing? how would you do it

OpenStudy (anonymous):

which derivatives do you have to take. the length and width?

OpenStudy (phi):

you start by writing down an equation that has area , length and width in it.

OpenStudy (phi):

I assume you know A= L*W

OpenStudy (anonymous):

yup

OpenStudy (phi):

in your problem, all three A, L and W are changing, so we treat all three as a variable step 1: write down the the equation A = L W step 2: take the derivative with respect to time dA/dt = d/dt(L W) on the right side, that is the derivative of two variables... we use the product rule d(u v) = u dv + v du

OpenStudy (phi):

can you take a stab at it? dA/dt = d/dt(L W)

OpenStudy (anonymous):

what would d/dt be?

OpenStudy (phi):

d/dt means take the derivative with respect to t it is like d/dx except we have t instead of x we use the normal rules: d/dt of t = dt/dt = 1 d/dt (y) = dy/dt d/dt (x) = dx/dt d/dt t^2 = 2t dt/dt = 2t follow ?

OpenStudy (anonymous):

ok so what are we taking the derivative of?

OpenStudy (phi):

dA/dt = d/dt(L W) do the right-hand side. L and W are variables (just like x and y)

OpenStudy (anonymous):

so it would be L(dW/dt)+(dL/dt)W?

OpenStudy (phi):

yes. looks good. As an aside, In physics people would write \( \dot{W} \) rather than dW/dt (Newton used this version) because it is easier. But it isn't easier to type. sometimes people use W' for dW/dt but it is never obvious that it is dW/dt (as opposed to dW/dx for example)

OpenStudy (anonymous):

i havent done physics yet. only in ninth grade

OpenStudy (phi):

you started with A = L W took the derivative with respect to time and found dA/dt = L dW/dt + W dL/dt now if we know the stuff on the right hand side, we can compute dA/dt

OpenStudy (phi):

The problem gives all the info we need

OpenStudy (anonymous):

L=30 W=25 dW/dt=5 dL/dt=8

OpenStudy (anonymous):

150+200=350?

OpenStudy (anonymous):

yup i got it. thanks so much. you are an amazing helper

OpenStudy (phi):

yes. But it is a good habit to keep track of the units. dA/dt = 350 cm^2/sec

OpenStudy (anonymous):

ok thanks so much!!!!!!